OpenStudy (anonymous):

Find S20 for the arithmetic series 8 + 16 + 24 + 32 +...

OpenStudy (anonymous):

160?

OpenStudy (anonymous):

\[S _{n} = 8+8(n-1)\] since your n is 20, subsitute that into the equation and you get \[S _{20} = 8+8(20-1)\] = 8 + 8(19) =160

OpenStudy (anonymous):

thankssss :)